OpenStudy (anonymous):

HEELP PLEASE!! A pushes horizontally eastward on a 1.2 kg textbook initially resting 1.3m from the edge of a long table. The coefficient of static friction is 0.24 and the coefficient of kinetic friction is 0.18. The force is just large enough to get the book to start moving, and is maintained for 1.0s after the motion has begun. Determine the acceleration of the book while the student is applying the push.

OpenStudy (anonymous):

The answer is 0.59 m/s^2 [E]

OpenStudy (anonymous):

1.2 kg weights (1.2)(9.8 N/kg) = 11.76 N To start it moving, static friction must be overcome with F = force = (0.24)(11.76N) = 2.82N Once it is moving, it resists with frictional force (0.18)(11.76N)=2.12N Leaving f=2.82 - 2.12 = 0.7N to accelerate the mass. acceleration = m/f = 1.2 kg/0.7N = 1.7 m/s^2